OpenStudy (anonymous):

The probability of winning a certain prize is 5% What are the odds of winning the prize?

OpenStudy (anonymous):

Please help

OpenStudy (tkhunny):

p(winning) = 5% = 0.05 p(losing) = 1 - p(winning) = 1 - 5% = 95% = 0.95 Odds of Winning: 0.05:0.95 Odds of Losing: 0.95:0.05 Odds and percentages are not quite the same thing.

OpenStudy (anonymous):

Thx